INAF queue observing, 2020 Dec 22 - Dec 23

Observer: A. Cardwell
Telescope Operator: Davide Huerta
Astronomers from Italy: F. Cusano, A. Rossi
Instrument: PESI-BINO

Summary

Due to the covid-19 pandemic situation these observations have been performed by LBTO personel
through the indications of LBT Italian observers (A. Rossi and F. Cusano) connected from Italy.

Wind gust above the limits at the beginning of the night we stay closed at sunset, we opened up at 4:50. Poor to awful seeing all night, and frequent wind gusts over the shutdown limit.
Ultimately we decided to spend the whole night with PEPSI, and did take a few hours of good data.

Night duration (12 degree el twilight) 01:20 - 13:20 UT = 12 h 00m
Observing time: 4h 0min ( 34%)
Weather time loss: 8h 0min ( 66%)
technical time loss h min ( %)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Clementini 3026, 1107 PEPSI-BINO        
Lanzafame 2351616 PEPSI-BINO        

Details

Detailed log of the nigh by A. Cardwell at this link: https://drive.google.com/file/d/1Nx8FY2XmLWK9oqNFcLvwn71aqp8llWI8/view?usp=sharing

INAF queue observing, 2020 Dec 23 - Dec 24

Observer: A. Cardwell
Telescope Operator: Davide Huerta
Astronomers from Italy: F. Cusano, A. Rossi
Instrument: LUCI-BINO

Summary

Due to the covid-19 pandemic situation these observations have been performed by LBTO personel
through the indications of LBT Italian observers (A. Rossi and F. Cusano) connected from Italy.

Condition are good to open at the beginning of the night, but internet connection was lost to and from the mountain a 10 mins after 12 degree twilight. It was not recovered during the night.
Using the 4G hotspot we talked the OSA through the LUCI observations for the night. Progress was slow, but we did what we could. Average seeing was between 2” and 3” for most of the night, improving to 1” to 1.5” for the last hour or two of the night.

Night duration (12 degree el twilight) 01:20 - 13:20 UT = 12 h 00m
Observing time: 3h min ( 25%)
Weather time loss: 0h min ( %)
technical time loss 9h min ( 75%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Di Criscienzo Peg LUCI-BINO        
Rossi GRB201216C LUCI-BINO H-ima     seeing>2, not good
Rossi GRB201221D LUCI-BINO J-K ima 0.6h    
Antoniucci Fu Ori LUCI-BINO       completed? check quality

Details

Detailed log of the nigh by A. Cardwell at this link: https://drive.google.com/file/d/18f_LPHL2kTXSMIn7Q9yPFvSthzHsLYbV/view?usp=sharing

INAF queue observing, 2020 Dec 24 - Dec 25

Observer: A. Cardwell
Telescope Operator: Davide Huerta
Astronomers from Italy: F. Cusano, A. Rossi
Instrument: MODS-BINO

Summary

Due to the covid-19 pandemic situation these observations have been performed by LBTO personel
through the indications of LBT Italian observers (A. Rossi and F. Cusano) connected from Italy.

At the sunset the sky is cloudy. DOME CLOSED. We opened at 7:20 UT

Night duration (12 degree el twilight) 01:20 - 13:20 UT = 12 h 00m
Observing time: 4h 30min ( 38%)
Weather time loss: 7h 30min ( 62%)
technical time loss h min ( %)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
STD GD71 MODS1+2 red-grating     std for Moretti
Moretti_DDT PSOJ0728+33 MODS1+2 red-grating 1"slit 1200*12=4h   completed

Details

Detailed log of the nigh by A. Cardwell at this link: https://drive.google.com/file/d/1c3_AzcXASCUmVyzA1q7DkwpM1DQDmxt1/view?usp=sharing

INAF queue observing, 2020 Dec 25 - Dec 26

Observer: A. Cardwell
Telescope Operator: Davide Huerta
Astronomers from Italy: F. Cusano, A. Rossi
Instrument: LUCI-BINO

Summary

Due to the covid-19 pandemic situation these observations have been performed by LBTO personel
through the indications of LBT Italian observers (A. Rossi and F. Cusano) connected from Italy.
At the sunset the sky is clear but humidity is high. We wait for humidity to go down.

Night duration (12 degree el twilight) 01:20 - 13:20 UT = 12 h 00m
Observing time: 11h min ( 92%)
Weather time loss: h min ( %)
technical time loss 1h min ( 8%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Davanzo_DDT GRB200826A LUCI1 AO, imaging, Hband 45min/1.6h   BIN3; low image quality
telluric   LUCI1+2 MOS     for Gargiulo
Gargiulo VAN001 LUCI1+2 MOS 1h+1h    
telluric HIP39540 LUCI1+2       for Nardini
Nardini_4 J0945 LUCI1+2   1.67h   took a few more since we started at lower airmass than recested completed
Nardini_4 J1014 LUCI1+2   0.2/0.1h   took double time to have all offsets completed
telluric HIP39540 LUCI1+2       for Nardini
Nardini_4 J0947 LUCI1+2   0.23h   completed
Nardini_4 J1027 LUCI1+2   0.2/0.1h   took double time to have all offsets completed
telluric HIP52849 LUCI1+2       for Nardini
telluric HIP43018 LUCI1+2       for Saxena
Saxena_39 COSMOS1 LUCI1+2 MOS 3.5/15h   seeing degraded in last 30min; Remove star#2 to allign

Details

Detailed log of the nigh by A. Cardwell at this link: https://drive.google.com/file/d/1bDGyuEmTzVlqtNe-HTbTSMqcLEwL-eH0/view?usp=sharing

-- AndreaRossi - 26 Dec 2020
Topic revision: r3 - 07 Jan 2021, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback